# Communication-Efficient Federated Fine-Tuning of Language Models via Dynamic Update Schedules

### v1

@@ Abstract.
Federated learning (FL) makes it possible to train models on data that would otherwise remain untapped and inaccessible. Simultaneously, pre-trained language models (LMs) have emerged as indispensable tools in modern workflows. These models exhibit extraordinary capabilities and are easily adapted to downstream tasks. This opens one of the most exciting frontiers in FL: fine-tuning LMs. However, a persistent challenge in FL is the frequent, rigid communication of parameters, a problem which is magnified by the sheer size of these modern models. Currently, the FedOpt family of algorithms is the prevailing approach in FL, though it relies on fixed, heuristic intervals for model synchronization. Recently, the FDA algorithm introduced a dynamic alternative by monitoring training progress, but it came with its own drawbacks; namely, a hard-to-tune threshold parameter and a rigid synchronization scheme. In this work, we introduce the FDA-Opt family of algorithms—a unified generalization that extends the principles behind both FDA and FedOpt, while resolving their core limitations. We evaluate our approach on fine-tuning LMs across a range of downstream NLP tasks, and demonstrate that it consistently outperforms FedOpt—even when FDA-Opt operates under hyper-parameter settings originally optimized for its competitors. In other words, we show that FDA-Opt is a practical, drop-in replacement for FedOpt in modern FL libraries and systems: it requires no additional configuration and delivers superior performance out of the box.
Federated Learning, Deep Learning, Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Machine learningccs: Computing methodologies Distributed algorithms
1. INTRODUCTION
@@
In today’s data-driven landscape, leveraging data effectively has become both an opportunity and a challenge. Sensitive data, such as patient records in healthcare or personal information from mobile applications, is often siloed due to privacy concerns (Yan et al., 2024) and stringent regulations (Voigt and Bussche, 2017). While necessary for protecting user rights, these restrictions hinder the potential of machine learning solutions, as they prevent data from being centralized for training. Federated Learning (FL) (Kairouz et al., 2021; McMahan et al., 2017) emerges as a transformative paradigm by enabling models to be trained directly on the devices or organizations that own the data, bypassing restrictions and privacy concerns.
@@
At the same time, pre-trained Language Models (LMs) have demonstrated extraordinary generalization power that enables them to be readily adapted—through techniques like fine-tuning—to a broad range of downstream tasks (Raffel et al., 2020). By following the FL paradigm, these powerful models can be trained on vast, previously untapped data sources, enabling new breakthroughs and applications in natural language processing (NLP) (Woisetschläger et al., 2024; Wu et al., 2024). However, FL introduces a fundamental challenge: the frequent exchange of model updates. This issue is only exacerbated by the immense size of modern transformer-based models, making communication overhead a critical bottleneck (Woisetschläger et al., 2024).
@@
Figure 1. A Federated Learning Round
The widely adopted answer to this question is the FedOpt family of algorithms (Reddi et al., 2021), which prescribe ending rounds after a fixed number of local steps or epochs. In this approach, each client processes its dataset one or more times before sending model updates back to the server. However, this method has two significant limitations. First, it assumes the training process is static, with configurations predefined throughout training, without taking into account real-time training dynamics. Second, the choice of the number of local steps is arbitrary and lacks justification. Lower values may improve convergence but skyrocket communication overhead, while higher values reduce communication cost but risk unpredictable training behavior or, in many cases, non-convergence (Yu et al., 2019).
Recently, this trade-off has motivated the development of a dynamic approach to round termination: Federated Dynamic Averaging (FDA(Theologitis et al., 2025). Unlike static methods, FDA monitors the training state in real-time and adjusts the duration of each round accordingly. Specifically, FDA tracks the model variance—a gauge of whether training is progressing well or poorly—and ensures it remains below a prespecified threshold. As long as the variance is within this threshold, local training continues. When the variance exceeds the threshold, the round is terminated, and the server collects the models. This adaptive approach has been shown to significantly reduce communication overhead, achieving orders-of-magnitude improvements in communication compared to FedOpt algorithms.
However, the original formulation of FDA (Theologitis et al., 2025) has certain limitations. First, it enforces a rigid synchronization scheme, requiring clients to exchange information with the server synchronously after every learning step, which can bottleneck a real-world FL system. Second, while FDA successfully eliminates the need to tune the number of local steps, it introduces a new parameter—the variance threshold—which is highly nontrivial to calibrate. In addition to these limitations, the experimental comparisons between FDA and FedOpt were between best-case configurations for each algorithm, with different optimizers and hyper-parameters, raising the question whether FDA consistently outperforms FedOpt in an apples-to-apples scenario.
@@
The remainder of this paper is organized as follows. Section 2 introduces the foundational concepts of Federated Learning (FL), including the FedOpt family of algorithms and the dynamic variance-based FDA approach. Section 3 presents our proposed FDA-Opt algorithm. Section 4 outlines the experimental approach. Section 5 provides a detailed analysis of experimental results. Section 6 reviews related work. Finally, Section 7 summarizes the main findings of this work.
2. PRELIMINARIES
Before diving into our proposed algorithm, it is essential to understand the key concepts and methods that form its foundation. In this section, we break down the FL training process, explore the widely adopted FedOpt family of algorithms, and introduce the recently proposed dynamic FDA approach. Together, these elements set the stage for understanding our subsequent contributions and the idea behind our generalized FDA-Opt algorithm in Section 3.
2.1. Federated Learning Basics
@@
Instead of centralizing data on a powerful cluster for training, FL reverses the traditional approach: training and computation are brought to the data, that is, to the clients. Each client trains a model locally and shares updates with a central server, which consolidates them into a common global model, ${\bm{w}}\in\mathbb{R}^{d}$. The collective goal is to minimize the overall training loss across all clients, which can be expressed as the following distributed optimization problem (Woisetschläger et al., 2024):
(1) $\underset{{\bm{w}}\in{\mathbb{R}}^{d}}{\mathrm{minimize}}\;\;F({\bm{w}})\triangleq\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}F_{k}({\bm{w}})$
Here, $F_{k}({\bm{w}})\triangleq\mathbb{E}_{\zeta_{k}\sim\mathcal{D}_{k}}\left[\ell({\bm{w}};\zeta_{k})\right]$ is the local objective function for client $k$, and $\ell({\bm{w}};\zeta_{k})$ denotes the loss for a data sample $\zeta_{k}$ given the model ${\bm{w}}$.
2.2. Generalized Federated Averaging
The most widely used algorithm for solving (1) is Federated Averaging (FedAvg(McMahan et al., 2017). This method divides the FL training process into iterative rounds. At the beginning of round $t$, the server broadcasts the current global model ${\bm{w}}_{t}$ to a selected cohort of participating clients $\mathcal{S}_{t}\subseteq\mathcal{K}$. Each client $k\in\mathcal{S}_{t}$ initializes its local model as ${\bm{w}}^{(k)}_{t,0}={\bm{w}}_{t}$, and performs $\tau$ local updates of Stochastic Gradient Descent (SGD) using its private dataset $\mathcal{D}_{k}$. After training, the client sends its updated model ${\bm{w}}^{(k)}_{t,\tau}$ back to the server. The server then aggregates the updates—typically via averaging [^1] To be precise, the aggregation is a weighted average based on the number of samples $|\mathcal{D}_{k}|$ owned by each client; however, to simplify notation, and because the extension to a weighted average is straightforward, we will use averaging throughout this paper.—to form the new global model, ${\bm{w}}_{t+1}$.
An indirect yet highly effective way to mitigate the communication burden of FL is to accelerate convergence which led to efforts to incorporate accelerated optimization techniques—such as Adam (Kingma and Ba, 2015)—into the FL setting. However, FL clients typically lack enough data for robust statistical coverage and are often stateless, making adaptive optimization at the client level impractical. Instead, the key idea is to shift adaptive optimization to the server.
@@ 2.2. Generalized Federated Averaging
which we refer to as the “pseudo-gradient”. The term “pseudo” is used because even though ${\bm{g}}_{t,\tau}$ is not a traditional gradient—it was not derived from a loss function via backpropagation—it can still be treated in the same way (Reddi et al., 2021). The server then applies a server-side optimizer, ServerOpt, to update the global model using this “pseudo”-gradient. Importantly, since ${\bm{g}}_{t,\tau}$ aggregates updates from multiple clients, it offers sufficient and trustworthy statistics, and the server can easily maintain the optimizer state across rounds. In effect, this addresses the two challenges that make adaptive methods impractical on clients: unreliable statistics, and statelessness.
This approach underpins the FedOpt family of algorithms (Reddi et al., 2021), which correspond to the non-highlighted lines of Algorithm 1 in Section 3, alongside our proposed FDA-Opt algorithm (highlighted lines). We present them together, as FDA-Opt generalizes the core functionality of FedOpt. Additionally, Table 1 summarizes the FedOpt family. Notably, FedAvg—which simply averages updates—is mathematically equivalent to SGD with a learning rate of $1.0$.
Table 1. Overview of the FedOpt and FDA-Opt families of algorithms. Depending on ClientOpt and ServerOpt a different name is derived for the FL algorithm. FDA-Opt is our proposed algorithm, which will be introduced in Section 3.
@@
FedAvg (McMahan et al., 2017) FDA-SGD SGD SGD (lr $=1.0$)
FedAvgM (Hsu et al., 2019) FDA-SGDM SGD SGDM (Sutskever et al., 2013)
FedAdam (Reddi et al., 2021) FDA-Adam SGD Adam (Kingma and Ba, 2015)
FedAdamW (Woisetschläger et al., 2024) FDA-AdamW SGD AdamW (Loshchilov and Hutter, 2019)
FedAdaGrad (Reddi et al., 2021) FDA-AdaGrad SGD AdaGrad (Duchi et al., 2011)
2.3. Federated Dynamic Averaging
Federated Dynamic Averaging (FDA(Theologitis et al., 2025) is another algorithm for solving (1), with a key distinction from FedAvg: instead of fixing the number of local steps $\tau$ per round, FDA dynamically decides when clients should stop training and return their models. Specifically, it monitors a metric called model variance, and terminates local training once this variance exceeds a predefined threshold $\Theta$.
@@ 2.3. Federated Dynamic Averaging
Here, $H$ is a function that maps the global state to a variance approximation (we refer to the original work for the full details (Theologitis et al., 2025)). Importantly, if a threshold violation is detected, $\nu_{t,i}>\Theta$, training stops, and the server collects models; otherwise, local training proceeds to the next SGD step, $i+1$, and Steps 2.1–2.4 repeat (Figure 2).
@@
Figure 2. Mechanics of local training, in round $t$, under FDA
3. FDA-Opt: GENERALIZED FEDERATED DYNAMIC AVERAGING
@@
3.1. Optimizers
In the original FDA algorithm, the server-side aggregation relied on a simple averaging scheme. However, extensive empirical evidence highlights the effectiveness of adaptive optimizers in FL (Reddi et al., 2021). Naturally, we extend the server-side aggregation step by introducing the use of arbitrary optimizers in FDA-Opt, denoted as ServerOpt.
@@ 3.2. Model Variance
${\bm{W}}_{t,i}\triangleq\left[{\bm{w}}_{t,i}^{(k_{1})},{\bm{w}}_{t,i}^{(k_{2})},\ldots,{\bm{w}}_{t,i}^{(k_{N})}\right]$
@@
Additionally, as detailed in Section 2.2, each client’s model change, ${\bm{\Delta}}^{(k)}_{t,i}$, captures the drift after the $i$-th local training step. Furthermore, their average, ${\bm{g}}_{t,i}$, can be interpreted as a “pseudo”-gradient:
@@
Definition. The model variance quantifies the spread (or dispersion) of the client models around their average—indicating how compactly they are clustered in parameter space. At round $t$, after $i$ local training steps, it can be written as (Theologitis et al., 2025):
@@
Figure 3. Variance progression during the first $100$ rounds of training with FedOpt using RoBERTa on the MRPC dataset
@@
3.3. The FDA-Opt Algorithm
High-Level Overview. Having established model variance as a key indicator of training dynamics—with distinct behaviors and trends—we now leverage these insights for a unique goal: to extend local training duration beyond traditional limits without compromising convergence. To this end, we the propose FDA-Opt, presented in Algorithm 1. The highlighted lines represent the core additions that distinguish FDA-Opt: namely, the logic for monitoring model variance and dynamically terminating rounds. These additions act as an adapter to the original FedOpt framework (non-highlighted lines, also detailed in Section 2.2), effectively generalizing it with an dynamic round termination scheme.
@@ 3.3. The FDA-Opt Algorithm
From this point onward, the algorithm proceeds using the standard round-based approach described in Section 2.2. At the beginning of each round $t$, a cohort of clients $\mathcal{S}_{t}$ is sampled. These clients train locally for up to $\tilde{\tau}$ steps (Lines 3–9). At the end of training—whose actual length may vary—we compute the “pseudo”-gradient from model changes and update the global model accordingly (Lines 18 and 21). This process is repeated for a total of $T$ rounds. What sets FDA-Opt apart is its variance monitoring logic.
Specifically, during local training, clients periodically synchronize compact state information (Lines 11–12) to approximate the model variance (Line 13). However, querying variance too frequently (e.g., after every local step) can completely bottleneck a real-world FL system. This is not due to communication volume (the local state states are small), but due to the synchronous nature of the operation. To mitigate this, we introduce a configurable set of step indices, $\mathcal{I}_{\text{query}}\subseteq[1,2,...,\tilde{\tau}]$, which determines the exact points where variance queries are performed (Line 10). In contrast to the original FDA formulation (Theologitis et al., 2025), which used $\mathcal{I}_{\text{query}}=[1,2,...,\tilde{\tau}]$ (i.e., checking after every step), our method leaves this set user-defined. In practice, we encourage it to be significantly sparser. As we will show later, in our setup the variance is queried only once per epoch.
Lastly, following the discussion on variance behavior in Section 3.2, we allow the threshold value, $\Theta_{t}$, to evolve across training rounds. Rather than using a fixed value, the threshold is updated dynamically based on the observed training dynamics. Specifically, we define a function ThresholdAdjust, which takes as input the current threshold $\Theta_{t}$, the actual variance value at the end of the round, and the final step index $s_{t}$ where the round terminated (Lines 19–20). Intuitively, these inputs capture the key information needed to answer two questions: (1) was the current threshold too high or too low? and (2) did the round terminate at a reasonable point (e.g., $s_{t}=1$ should raise concerns)? In the next section, we will explore practical choices for this function. Notably, computing the variance in Line 19 requires no additional work, as it reuses quantities—the drifts and “pseudo”-gradient—that are already available (Lines 17–18).
@@
1:Input: Initial ${\bm{w}}_{0}$; ClientOpt, ServerOpt; Total rounds $T$; Local training steps $\tau$, specifically tuned for FedOpt (Reddi et al., 2021)
@@
10:         Compute a gradient estimate ${\bm{g}}^{(k)}_{t,i-1}$ of $\nabla F_{k}({\bm{w}}^{(k)}_{t,i-1})$
@@
21:   $\nu ar_{t}=\mathrm{Var}({\bm{W}}_{t,s_{t}})$triangleright Compute actual variance using (3)
@@
3.4. Default Configuration for FDA-Opt
We adopt a conservative approach in the configurations below. These are the exact settings used in our experiments, as they yield stable and reliable performance across tasks. However, we believe that there is room for improvement, possibly through more aggressive choices (e.g., a larger linear coefficient of increase of $\tau$).

### v2

# FDA-Opt: Communication-Efficient Federated Fine-Tuning of Language Models

@@ Abstract
Federated Learning (FL) enables the utilization of vast, previously inaccessible data sources. At the same time, pre-trained Language Models (LMs) have taken the world by storm and for good reason. They exhibit remarkable emergent abilities and are readily adapted to downstream tasks. This opens one of the most exciting frontiers in FL: fine-tuning LMs. Yet, a persistent challenge in FL is the frequent, rigid communication of parameters—a problem magnified by the sheer size of these contemporary models. The FedOpt family of algorithms has become the go-to approach for FL, relying on fixed but arbitrary intervals for model exchanges. Recently, the FDA algorithm prescribed a dynamic approach by monitoring the training progress. However, it introduced a hard-to-calibrate parameter and imposed a rigid synchronization scheme. In this work, we address these limitations by proposing the FDA-Opt family of algorithms—a unified generalization of both FDA and FedOpt. Our experimental evaluation focuses on fine-tuning LMs on downstream NLP tasks and demonstrates that FDA-Opt outperforms FedOpt even when it is configured with hyper-parameters specifically optimized for the latter. In other words, we show that FDA-Opt is a practical, drop-in replacement for FedOpt in modern FL libraries and systems: it requires no additional configuration and delivers superior performance out of the box.
Index Terms:
Federated Learning, Deep Learning, Language Models.
I Introduction
@@
In today’s data-driven landscape, leveraging data effectively has become both an opportunity and a challenge. Sensitive data, such as patient records in healthcare or personal information from mobile applications, is often siloed due to privacy concerns [1, 2] and stringent regulations [3]. While necessary for protecting user rights, these restrictions hinder the potential of machine learning solutions, as they prevent data from being centralized for training. Federated Learning (FL) [4, 5] emerges as a transformative paradigm by enabling models to be trained directly on the devices or organizations that own the data, bypassing restrictions and privacy concerns.
@@
At the same time, pre-trained Language Models (LMs) have demonstrated extraordinary generalization power that enables them to be readily adapted—through techniques like fine-tuning—to a broad range of downstream tasks [6, 7]. By following the FL paradigm, these powerful models can be trained on vast, previously untapped data sources, enabling new breakthroughs and applications in natural language processing (NLP) [8, 9]. However, FL introduces a fundamental challenge: the frequent exchange of model updates. This issue is only exacerbated by the immense size of modern transformer-based models, making communication overhead a critical bottleneck [10].
@@
Figure 1: A Federated Learning Round
The widely adopted answer to this question is the FedOpt family of algorithms [11], which prescribe ending rounds after a fixed number of local steps or epochs. In this approach, each client processes its dataset one or more times before sending model updates back to the server. However, this method has two significant limitations. First, it assumes the training process is static, with configurations predefined throughout training, without taking into account real-time training dynamics. Second, the choice of the number of local steps is arbitrary and lacks justification. Lower values may improve convergence but skyrocket communication overhead, while higher values reduce communication cost but risk unpredictable training behavior or, in many cases, non-convergence [12].
Recently, this trade-off has motivated the development of a dynamic approach to round termination: Federated Dynamic Averaging (FDA[13]. Unlike static methods, FDA monitors the training state in real-time and adjusts the duration of each round accordingly. Specifically, FDA tracks the model variance—a gauge of whether training is progressing well or poorly—and ensures it remains below a prespecified threshold. As long as the variance is within this threshold, local training continues. When the variance exceeds the threshold, the round is terminated, and the server collects the models. This adaptive approach has been shown to significantly reduce communication overhead, achieving orders-of-magnitude improvements in communication compared to FedOpt algorithms.
However, the original formulation of FDA [13] has certain limitations. First, it enforces a rigid synchronization scheme, requiring clients to exchange information with the server synchronously after every learning step, which can bottleneck a real-world FL system. Second, while FDA successfully eliminates the need to tune the number of local steps, it introduces a new parameter—the variance threshold—which is highly nontrivial to calibrate. In addition to these limitations, the experimental comparisons between FDA and FedOpt were between best-case configurations for each algorithm, with different optimizers and hyper-parameters, raising the question whether FDA consistently outperforms FedOpt in an apples-to-apples scenario.
@@
The remainder of this paper is organized as follows. Section II introduces the foundational concepts of Federated Learning (FL), including the FedOpt family of algorithms and the dynamic variance-based FDA approach. Section III presents our proposed FDA-Opt algorithm. Section IV outlines the experimental approach. Section V provides a detailed analysis of experimental results. Section VI reviews related work. Finally, Section VII summarizes the main findings of this work.
II Preliminaries
Before diving into our proposed algorithm, it is essential to understand the key concepts and methods that form its foundation. In this section, we break down the FL training process, explore the widely adopted FedOpt family of algorithms, and introduce the recently proposed dynamic FDA approach. Together, these elements set the stage for understanding our subsequent contributions and the idea behind our generalized FDA-Opt algorithm in Section III.
II-A Federated Learning Basics
@@
Instead of centralizing data on a powerful cluster for training, FL reverses the traditional approach: training and computation are brought to the data, that is, to the clients. Each client trains a model locally and shares updates with a central server, which consolidates them into a common global model, ${\bm{w}}\in\mathbb{R}^{d}$. The collective goal is to minimize the overall training loss across all clients, which can be expressed as the following distributed optimization problem [10]:
$\underset{{\bm{w}}\in{\mathbb{R}}^{d}}{\mathrm{minimize}}\;\;F({\bm{w}})\overset{\Delta}{=}\frac{1}{|\mathcal{K}|}\sum_{k\in\mathcal{K}}F_{k}({\bm{w}})$ (1)
Here, $F_{k}({\bm{w}})\overset{\Delta}{=}\mathbb{E}_{\zeta_{k}\sim\mathcal{D}_{k}}\left[\ell({\bm{w}};\zeta_{k})\right]$ is the local objective function for client $k$, and $\ell({\bm{w}};\zeta_{k})$ denotes the loss for a data sample $\zeta_{k}$ given the model ${\bm{w}}$.
II-B Generalized Federated Averaging
The most widely used algorithm for solving (1) is Federated Averaging (FedAvg[5]. This method divides the FL training process into iterative rounds. At the beginning of round $t$, the server broadcasts the current global model ${\bm{w}}_{t}$ to a selected cohort of participating clients $\mathcal{S}_{t}\subseteq\mathcal{K}$. Each client $k\in\mathcal{S}_{t}$ initializes its local model as ${\bm{w}}^{(k)}_{t,0}={\bm{w}}_{t}$, and performs $\tau$ local updates of Stochastic Gradient Descent (SGD) using its private dataset $\mathcal{D}_{k}$. After training, the client sends its updated model ${\bm{w}}^{(k)}_{t,\tau}$ back to the server. The server then aggregates the updates—typically via averaging [^1] To be precise, the aggregation is a weighted average based on the number of samples $|\mathcal{D}_{k}|$ owned by each client; however, to simplify notation, and because the extension to a weighted average is straightforward, we will use averaging throughout this paper.—to form the new global model, ${\bm{w}}_{t+1}$.
An indirect yet highly effective way to mitigate the communication burden of FL is to accelerate convergence which led to efforts to incorporate accelerated optimization techniques—such as Adam [14]—into the FL setting. However, FL clients typically lack enough data for robust statistical coverage and are often stateless, making adaptive optimization at the client level impractical. Instead, the key idea is to shift adaptive optimization to the server.
@@ II-B Generalized Federated Averaging
which we refer to as the “pseudo-gradient”. The term “pseudo” is used because even though ${\bm{g}}_{t,\tau}$ is not a traditional gradient—it was not derived from a loss function via backpropagation—it can still be treated in the same way [11]. The server then applies a server-side optimizer, ServerOpt, to update the global model using this “pseudo”-gradient. Importantly, since ${\bm{g}}_{t,\tau}$ aggregates updates from multiple clients, it offers sufficient and trustworthy statistics, and the server can easily maintain the optimizer state across rounds. In effect, this addresses the two challenges that make adaptive methods impractical on clients: unreliable statistics, and statelessness.
This approach underpins the FedOpt family of algorithms [11], which correspond to the non-highlighted lines of Algorithm 1 in Section III, alongside our proposed FDA-Opt algorithm (highlighted lines). We present them together, as FDA-Opt generalizes the core functionality of FedOpt. Additionally, Table I summarizes the FedOpt family. Notably, FedAvg—which simply averages updates—is mathematically equivalent to SGD with a learning rate of $1.0$.
TABLE I: Overview of the FedOpt and FDA-Opt families of algorithms. Depending on ClientOpt and ServerOpt a different name is derived for the FL algorithm. FDA-Opt is our proposed algorithm, which will be introduced in Section III.
@@
FedAvg [5] FDA-SGD SGD SGD (lr $=1.0$)
FedAvgM [15] FDA-SGDM SGD SGDM [16]
FedAdam [11] FDA-Adam SGD Adam [14]
FedAdamW [8] FDA-AdamW SGD AdamW [17]
FedAdaGrad [11] FDA-AdaGrad SGD AdaGrad [18]
II-C Federated Dynamic Averaging
Federated Dynamic Averaging (FDA[13] is another algorithm for solving (1), with a key distinction from FedAvg: instead of fixing the number of local steps $\tau$ per round, FDA dynamically decides when clients should stop training and return their models. Specifically, it monitors a metric called model variance, and terminates local training once this variance exceeds a predefined threshold $\Theta$.
@@ II-C Federated Dynamic Averaging
Here, $H$ is a function that maps the global state to a variance approximation (we refer to the original work for the full details [13]). Importantly, if a threshold violation is detected, $\nu_{t,i}>\Theta$, training stops, and the server collects models; otherwise, local training proceeds to the next SGD step, $i+1$, and Steps 2.1–2.4 repeat (Figure 2).
@@
Figure 2: Mechanics of local training, in round $t$, under FDA
III FDA-Opt: Generalized Federated Dynamic Averaging
@@
III-A Optimizers
In the original FDA algorithm, the server-side aggregation relied on a simple averaging scheme. However, extensive empirical evidence highlights the effectiveness of adaptive optimizers in FL [11]. Naturally, we extend the server-side aggregation step by introducing the use of arbitrary optimizers in FDA-Opt, denoted as ServerOpt.
@@ III-B Model Variance
${\bm{W}}_{t,i}\overset{\Delta}{=}\left[{\bm{w}}_{t,i}^{(k_{1})},{\bm{w}}_{t,i}^{(k_{2})},\ldots,{\bm{w}}_{t,i}^{(k_{N})}\right]$
@@
Additionally, as detailed in Section II-B, each client’s model change, ${\bm{\Delta}}^{(k)}_{t,i}$, captures the drift after the $i$-th local training step. Furthermore, their average, ${\bm{g}}_{t,i}$, can be interpreted as a “pseudo”-gradient:
@@
Definition. The model variance quantifies the spread (or dispersion) of the client models around their average—indicating how compactly they are clustered in parameter space. At round $t$, after $i$ local training steps, it can be written as [13]:
@@
Figure 3: Variance progression during the first $100$ rounds of training with FedOpt using RoBERTa on the MRPC dataset
@@
III-C The FDA-Opt Algorithm
High-Level Overview. Having established model variance as a key indicator of training dynamics—with distinct behaviors and trends—we now leverage these insights for a unique goal: to extend local training duration beyond traditional limits without compromising convergence. To this end, we the propose FDA-Opt, presented in Algorithm 1. The highlighted lines represent the core additions that distinguish FDA-Opt: namely, the logic for monitoring model variance and dynamically terminating rounds. These additions act as an adapter to the original FedOpt framework (non-highlighted lines, also detailed in Section II-B), effectively generalizing it with an dynamic round termination scheme.
@@ III-C The FDA-Opt Algorithm
From this point onward, the algorithm proceeds using the standard round-based approach described in Section II-B. At the beginning of each round $t$, a cohort of clients $\mathcal{S}_{t}$ is sampled. These clients train locally for up to $\tilde{\tau}$ steps (Lines 3–9). At the end of training—whose actual length may vary—we compute the “pseudo”-gradient from model changes and update the global model accordingly (Lines 18 and 21). This process is repeated for a total of $T$ rounds. What sets FDA-Opt apart is its variance monitoring logic.
Specifically, during local training, clients periodically synchronize compact state information (Lines 11–12) to approximate the model variance (Line 13). However, querying variance too frequently (e.g., after every local step) can completely bottleneck a real-world FL system. This is not due to communication volume (the local state states are small), but due to the synchronous nature of the operation. To mitigate this, we introduce a configurable set of step indices, $\mathcal{I}_{\text{query}}\subseteq[1,2,...,\tilde{\tau}]$, which determines the exact points where variance queries are performed (Line 10). In contrast to the original FDA formulation [13], which used $\mathcal{I}_{\text{query}}=[1,2,...,\tilde{\tau}]$ (i.e., checking after every step), our method leaves this set user-defined. In practice, we encourage it to be significantly sparser. As we will show later, in our setup the variance is queried only once per epoch.
Lastly, following the discussion on variance behavior in Section III-B, we allow the threshold value, $\Theta_{t}$, to evolve across training rounds. Rather than using a fixed value, the threshold is updated dynamically based on the observed training dynamics. Specifically, we define a function ThresholdAdjust, which takes as input the current threshold $\Theta_{t}$, the actual variance value at the end of the round, and the final step index $s_{t}$ where the round terminated (Lines 19–20). Intuitively, these inputs capture the key information needed to answer two questions: (1) was the current threshold too high or too low? and (2) did the round terminate at a reasonable point (e.g., $s_{t}=1$ should raise concerns)? In the next section, we will explore practical choices for this function. Notably, computing the variance in Line 19 requires no additional work, as it reuses quantities—the drifts and “pseudo”-gradient—that are already available (Lines 17–18).
@@
1:Input: Initial ${\bm{w}}_{0}$; ClientOpt, ServerOpt; Total rounds $T$; Local training steps $\tau$, specifically tuned for FedOpt [11]
@@
10:         Compute gradient estimate ${\bm{g}}^{(k)}_{t,i-1}$ of $\nabla F_{k}({\bm{w}}^{(k)}_{t,i-1})$
@@
21:   $\nu ar_{t}=\mathrm{Var}({\bm{W}}_{t,s_{t}})$triangleright Compute actual var. using (3)
@@
III-D Default Configuration for FDA-Opt
We adopt a conservative approach in the configurations below. These are the exact settings used in our experiments, as they yield stable and reliable performance across all tasks. While we believe there is room for improvement—potentially through more aggressive configurations (e.g., a larger linear coefficient of increase of $\tau$)—we leave such tuning as a direction for future work.